# The Price of Uncertainty for Social Consensus

### v2

@@ Abstract
How hard is it to achieve consensus in a social network under uncertainty? In this paper we model this problem as a social graph of agents where each vertex is initially colored red or blue. The goal of the agents is to achieve consensus, which is when the colors of all agents align. Agents attempt to do this locally through steps in which an agent changes their color to the color of the majority of their neighbors. In real life, agents may not know exactly how many of their neighbors are red or blue, which introduces uncertainty into this process. Modeling uncertainty as perturbations of relative magnitude $1+\varepsilon$ to these color neighbor counts, we show that even small values of $\varepsilon$ greatly hinder the ability to achieve consensus in a social network. We prove theoretically tight upper and lower bounds on the price of uncertainty, a metric defined by Balcan et al. (2009) to quantify the effect of uncertainty in network games.
@@ 1 Introduction
In the past decade the rapid expansion of social media platforms has resulted in the formation of large and highly connected social networks. The evolution of these networks has been extensively studied in order to better understand how opinions held by agents are affected by their interactions with other people. In a social setting, individuals generally tend to modify their beliefs to more closely align with others that they frequently interact with (Moussaïd et al., 2013; Garimella et al., 2018). This generally causes opinions held by agents in a network to generally converge over time as interacting agents attempt to reach a consensus. However, previous research has shown that additional factors, such as a sparse network structure or agents who refuse to change their opinion, can hinder consensus formation. (Acemoğlu et al., 2013; Hegselmann and Krause, 2002; Moussaïd et al., 2013). The vast majority of social network models in the existing literature assume that agents have accurate information about the opinions held by others. However, in large networks where vertices generally have many neighbors and where interactions between agents are often brief, agents are often unable to make accurate judgments about the opinions of other agents such as the number of their consenting or dissenting peers (Xiong and Liu, 2014).
@@
Previous work modeling uncertainty in games has focused on the failure of agents to always pick the action with the optimal payoff. For example, in the quantal response model, agents sometimes deviate from optimal behavior with strategies providing higher payoffs being chosen more often than those providing lower payoffs (Penna, 2018; Blume, 1993; Alós-Ferrer and Netzer, 2010). Alternatively, agents may occasionally fail to participate in the game, which can sometimes increase the social cost (Meir et al., 2012; Li et al., 2017).
@@
In this paper, we instead focus on the effect of uncertainty in agents’ observations rather than their actions. We model a social network as a graph with an agent at each vertex. Individuals have a choice between two colors, red or blue, and the goal of each agent is to be the same color as the majority of their neighbors. This is called a consensus game (Christodoulou et al., 2012; Fabrikant et al., 2004). We consider a model of uncertainty first proposed by Balcan et al. (2009) where agents experience small, adversarial perturbations of magnitude $1+\varepsilon$ to their observations of their neighbors. These perturbations capture the difficulty of ascertaining accurate information in large social networks, as the absolute margin of error increases proportionally to the number of neighbors an agent has. Alternatively, this model can also capture the effect of biased measurements, where agents may distort observations to agree with their beliefs (confirmation bias) or place undue emphasis on new information (novelty bias) (Ngampruetikorn and Stephens, 2016; Mao et al., 2024).
@@
The goal of this paper is to study how uncertainty affects behavior in consensus games as agents make potentially suboptimal updates to their colors. Balcan et al. (2009) define a social cost function as the number of edges whose endpoints have different colors. They quantify the effect of uncertainty using a measure called the price of uncertainty (PoU) of consensus games, defined as the maximum multiplicative factor by which the social cost can increase from any starting state. Even extremely small perturbations can allow adversarial perturbations to snowball out of control and cause the game to move to states of much higher social cost (Balcan et al., 2011).
@@
Our main results are tight asymptotic upper and lower bounds on the PoU for consensus games. Letting $n$ denote the number of agents participating in the consensus game, the previously best known asymptotic bounds due to Balcan et al. (2011) were a lower bound of $\Omega(n^{2}\varepsilon^{3})$ and an upper bound of $O(n^{2}\varepsilon)$. In this paper, we extend their work by proving that the PoU for consensus games is exactly $\Theta(n^{2}\varepsilon^{2})$. We believe that our result also holds irrespective of agent update ordering (i.e., the adversary can only control the magnitude of the perturbations), and this a potential future direction for research.
@@ 2 The Model
We consider a consensus game, where each of $n$ agents are located at distinct vertices of a simple undirected graph with $n$ vertices (Christodoulou et al., 2012). In this paper, we will use the words agent and vertex interchangeably. Each agent can choose a color from $\{0,1\}$. Let us refer to two agents as neighbors iff there exists an edge between the two agents. An agent’s cost is defined as the number of their neighbors that pick a different color, which we call bad neighbors. We will also call edges whose endpoints have different colors bad edges. Finally, for any state $S$ of the game, we define a social cost function $\operatorname{cost}(S)$ as the sum of the costs of every agent in the game.
Starting from a ground state $S_{0}$ with social cost at least 1, let $S_{t}$ be a state achieved after $t$ steps under uncertain best-response dynamics. Here, the term uncertain here refers to the fact that an agent might miscount their cost by a multiplicative factor of at most $1+\varepsilon$ for an uncertainty factor $\varepsilon>0$ that can depend on $n$. Define an uncertain best response as a best response that strictly increases the social cost as a result of uncertainty. By definition, an uncertain best response must involve a color switch. Note that we can assume $\varepsilon=\Omega\left(\frac{1}{n}\right)$, otherwise an agent cannot miscount its number of bad neighbors and hence there is no uncertainty. The price of uncertainty (PoU) of consensus games is defined as
@@
This paper obtains a lower bound for the PoU of consensus games of $\Omega\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\tilde{\Omega}(n^{-1/2})$. Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$. This is achieved using an intricate construction that amplifies the “snowball” effect of uncertainty in Balcan et al. (2011). Next, we prove that the PoU for consensus games is also bounded from above by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega(\varepsilon^{-1/4})$ by analyzing a clever monovariant. These bounds coincide, so we conclude that the asymptotically tight PoU for consensus games is $\Theta(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega(n^{-1/4})$, resolving an open problem of Balcan et al. (2011).
@@ 2.1 Main Technical Challenges
Lower Bound. In order to achieve a lower bound of $\Omega(\varepsilon^{3}n^{2})$, Balcan et al. (2011) utilizes an initializer gadget to set up initial colorings for an output component. This component consists of successive layers of vertices, where each layer has more vertices than the previous layer by a factor of $(1+\varepsilon)$, and every vertex is connected to all vertices on adjacent layers. This output component is responsible for generating $\Omega(\varepsilon n^{2})$ bad edges, and the initializer gadget requires $\Theta(1/\varepsilon^{2})$ starting bad edges, so this achieves a lower bound of $\Omega(\varepsilon^{3}n^{2})$. We have found that it is very difficult to increase the final number of bad edges produced by such an output for any construction because the graph immediately becomes too connected for enough agents to play uncertain best responses. Rather, we present an alternative construction for an initializer gadget in Section 3.1 that only requires $\Theta(1/\varepsilon)$ starting bad edges, which results in a construction for which the number of bad edges can increase by a factor of $O(\varepsilon^{2}n^{2})$.
Upper Bound. Generally, the goal of an adversary that wants to maximally blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best-response process. Consider a time in which there are $m^{2}$ bad edges. The most “compact” way to store these bad edges, that is, uses the least number of vertices, is a bipartite graph $K_{m,m}$. If each of the $m$ vertices on one side of the bipartite graph plays an uncertain best-response, we can blow up the number of bad edges by a multiplicative factor of $(1+\varepsilon)$. However, this requires $(1+\varepsilon)m$ additional new vertices in order to “absorb” the new bad edges produced by the uncertain best-response. Hence to increase the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
For example, following this heuristic for constant $\varepsilon$, that is $\varepsilon$ that is not a function of $n$, the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$. Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain bad response, can be turned bad again. Assuming most edges can eventually be turned bad again as is the case in our lower bound construction, the total number of bad edges at the end is
@@
Since there must be $\Omega(1/\varepsilon)$ bad edges at the start, we estimate the upper bound on price of uncertainty as $O(\varepsilon^{2}n^{2})$.
It appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain bad responses. The hope is that we can prove an upper bound on the growth rate in line with our estimate of $O(\varepsilon^{2}n^{2})$. Unfortunately, this approach fails, as illustrated by the following example.
@@
Consider a white vertex $V$ connected to $m$ white vertices and $m+1$ red vertices for $m=\Omega(1/\varepsilon)$. The $m$ white vertices are all connected to another white vertex $W$ and the $m+1$ red vertices are all connected to another red vertex $R$. We claim that we can have $V$ play $2m+1$ best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices:
@@
  • Let $V$ play an uncertain best response that switches its color to red, resulting in $m+1$ bad edges.
@@
  • Take the endpoint of one such bad edge adjacent to $W$ and switch its color so that there are now $m$ bad edges incident to $V$ and 1 bad edge incident to $W$.
@@
  • Switch the color of $V$ again, resulting in $m+1$ bad edges. Take the endpoint of one such bad edge adjacent to $R$ and switch its color so that there are now $m$ bad edges incident to $V$, 1 bad edge incident to $W$, and 1 bad edge incident to $R$.
@@
  • Repeat the above, at each step switching the color of a neighbor of $V$ whose color has not been switched before, until all neighbors of $V$ have had their colors switched once. There are now $m$ bad edges incident to $R$ and $m+1$ bad edges incident to $W$.
@@
Note that Example 1 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is white, $m$ of the vertices in the $(2m+1)$-size part is red, and $m+1$ of the vertices in the $(2m+1)$-size part is white to also roughly double the number of bad edges. The bad edge multiplicative growth rate of this process operating on $\Theta(m)$ vertices is 2, whereas in our initial estimate it was $1+\varepsilon$. Hence using this analysis it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step. We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ bad edges at the last stage using this strategy.
@@
Instead, one clever idea is that we should focus not on the total number of bad edges at every step but rather consider the step at which each vertex first plays an uncertain best response. This imposes a sequential order on the vertices from the first vertex to play an uncertain best response to the last. Recall that in order for a vertex with high degree to be able to play an uncertain best response, it must be the case that a large number of edges adjacent to this vertex must be bad. If we are to end up with $\Omega(\varepsilon n^{2})$ bad edges at the end, it must be the case that the later vertices in this sequence must have high degree on average. This requires a large number of existing bad edges to be generated from the vertices before the final vertices in the sequence, which in turn requires bad edges from the vertices before them, and so on.
@@
Our proof of the $O(\varepsilon^{2}n^{2})$ upper bound analyzes the degrees of vertices in this sequence and shows that the sequence of degrees cannot increase “too fast” if vertices can only play $(1+\varepsilon)$-uncertain best responses, where we realize the notion of “too fast” as a clever monovariant.
@@ 3 Lower Bound
We first improve the lower bound for the PoU for consensus games. Furthermore, we do so in the expanded domain $\varepsilon=\tilde{\Omega}(n^{-1/2})$. Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
@@ 3.1 Proof of Improved Lower Bound
One general idea is to construct a graph as a sequence of levels where each level contains slightly more vertices than the previous level. Starting from the first level, there are enough bad edges for all vertices in the first level to play uncertain best responses, which increases the number of bad edges by a multiplicative factor of $1+\varepsilon$. The fact that the vertices in the first level switched color implies that the vertices in the second level now have enough bad neighbors to play an uncertain best response, and this effect snowballs across the levels.
@@
We now make the above idea rigorous. we define a boosting operation between layers of the following form. Let $\ell>1$ and suppose that level $\ell-1$ has $\frac{1}{1+\frac{\varepsilon}{2}}\cdot m$ vertices for some $m$, level $\ell$ has $m$ vertices, and level $\ell+1$ has $\left(1+\frac{\varepsilon}{2}\right)\cdot m$ vertices. Suppose that every vertex in level $\ell$ is connected to every vertex in level $\ell-1$ and level $\ell+1$, so the degree of every vertex in level $\ell$ is $d_{\ell}=\left(\frac{1}{1+\frac{\varepsilon}{2}}+(1+\frac{\varepsilon}{2})\right)\cdot m$. If the vertices in levels $\ell$ and $\ell+1$ are all white and the vertices in level $\ell-1$ are all red, then the edges between levels $\ell-1$ and $\ell$ are all bad. Since $(1+\varepsilon)\cdot\frac{1}{1+\frac{\varepsilon}{2}}\geq\frac{d_{\ell}}{2}$ for all $\varepsilon\leq 1$ then all vertices in level $\ell$ are allowed to play uncertain best responses in sequence, turning level $\ell$ all red. Note now that the number of bad edges between levels $\ell$ and $\ell+1$ is $\left(1+\frac{\varepsilon}{2}\right)\cdot m^{2}$, which is a $1+O(\varepsilon)$ multiplicative increase from the number of original bad edges $\frac{1}{1+\frac{\varepsilon}{2}}\cdot m^{2}$ from before this boosting operation.
@@
We now present the construction. In order to implement the idea above, we must first find some way to create the first two layers of size $\frac{1}{\varepsilon}$ since a vertex must have degree at least $\frac{1}{\varepsilon}$ in order to play an uncertain best response. The initializer layers will have numbers of vertices
$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,2,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{1}{\varepsilon},\frac{1}{\varepsilon}$
@@
where the odd indexed layers cycle the consecutive integers from $\frac{1}{\varepsilon}$ to $\frac{2}{\varepsilon}$ and the numbers of vertices in even indexed layers increase by a factor of 2 whenever the odd indexed layer’s cycle repeats until the number reaches $\frac{1}{\varepsilon}$. Each vertex is connected to all vertices in the layer immediately after it except for vertices in odd indexed layers with $\frac{2}{\varepsilon}$ vertices, which are connected to the next layer to form a regular bipartite graph with the next layer where the vertices in the next layer are each connected to $\frac{1}{\varepsilon}$ vertices in the current layer. Note that the total number of initializer layers is $\frac{1}{\varepsilon}\cdot\log_{2}\frac{1}{\varepsilon}$ and each layer has at most $\frac{2}{\varepsilon}$ vertices, so the total number of vertices in the initializer layers is $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$. The initializer sequence has been constructed in such a way that that starting from a state in which the first layer is red and all other layers are white, the vertices in each layer can play uncertain best responses to change their color to red in sequence so that the last two layers form a $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ where the left side is red and the right half is white. Finally, we construct $\frac{1}{\varepsilon}$ white vertices in a parallel layer to the red left side of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ and connect each of them to every vertex in the right half of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$, so that each vertex in the right half now has $\frac{1}{\varepsilon}$ red neighbors and $\frac{1}{\varepsilon}$ white neighbors on the left. Note that no new bad edges are added. We will now leave all layers before the current layer frozen, that is, no layer in a layer before the current layer will ever change color again.
@@
After the current layer, construct a secondary set of layers with numbers of vertices
$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,2,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{2}{\varepsilon},\frac{2}{\varepsilon}.$

### v3

@@ Abstract
How hard is it to achieve consensus in a social network under uncertainty? In this paper we model this problem as a social graph of agents where each vertex is initially colored red or blue. The goal of the agents is to achieve consensus, which is when the colors of all agents align. Agents attempt to do this locally through steps in which an agent changes their color to the color of the majority of their neighbors. In real life, agents may not know exactly how many of their neighbors are red or blue, which introduces uncertainty into this process. Modeling uncertainty as perturbations of relative magnitude $1+\varepsilon$ to these color neighbor counts, we show that even small values of $\varepsilon$ greatly hinder the ability to achieve consensus in a social network. We prove theoretically tight upper and lower bounds on the price of uncertainty, a metric defined in previous work by Balcan et al. to quantify the effect of uncertainty in network games.
@@ 1 Introduction
In the past decade, the rapid expansion of social media platforms has resulted in the formation of large and highly connected social networks. The evolution of these networks has been extensively studied in order to better understand how opinions held by agents are affected by their interactions with other people. In a social setting, individuals often modify their beliefs to more closely align with others that they frequently interact with (Moussaïd et al., 2013; Garimella et al., 2018). This generally causes opinions held by agents in a network to converge over time as interacting agents attempt to reach a consensus. However, previous research has shown that additional factors, such as a sparse network structure or agents who refuse to change their opinions, can hinder consensus formation (Acemoğlu et al., 2013; Hegselmann and Krause, 2002; Moussaïd et al., 2013). The vast majority of social network models in the existing literature assume that agents have accurate information about the opinions held by others. However, in large networks where vertices generally have many neighbors and where interactions between agents are often brief, agents are often unable to make accurate judgments about the opinions of other agents, such as the number of their consenting or dissenting peers (Xiong and Liu, 2014).
@@
Previous work modeling uncertainty in games has focused on the failure of agents to always pick the action with the optimal payoff. For example, in the quantal response model, agents sometimes deviate from optimal behavior, with strategies providing higher payoffs being chosen more often than those providing lower payoffs (Penna, 2018; Blume, 1993; Alós-Ferrer and Netzer, 2010). Alternatively, agents may occasionally fail to participate in the game, which can sometimes increase the social cost (Meir et al., 2012; Li et al., 2017).
@@
In this paper, we instead focus on the effect of uncertainty in agents’ observations rather than their actions. We model a social network as a graph with an agent at each vertex. Individuals have a choice between two colors, red or blue, and the goal of each agent is to be the same color as the majority of their neighbors. This is called a consensus game (Christodoulou et al., 2012; Fabrikant et al., 2004). We consider a model of uncertainty first proposed by Balcan et al. (2009), where agents experience small, adversarial perturbations of relative magnitude $1+\varepsilon$ to their observations of their neighbors. These perturbations capture the difficulty of ascertaining accurate information in large social networks, as the absolute margin of error increases proportionally to the number of neighbors an agent has. Alternatively, this model can also capture the effect of biased measurements, where agents may distort observations to agree with their beliefs (confirmation bias) or place undue emphasis on new information (novelty bias) (Ngampruetikorn and Stephens, 2016; Mao et al., 2024).
@@
The goal of this paper is to study how uncertainty affects behavior in consensus games as agents make potentially suboptimal updates to their colors. Balcan et al. (2009) define a social cost function as the number of edges whose endpoints have different colors. They quantify the effect of uncertainty using a measure called the price of uncertainty (PoU) of consensus games, defined as the maximum multiplicative factor by which the social cost can increase from any starting state. Even extremely small adversarial perturbations can snowball out of control and cause the game to move to states of much higher social cost (Balcan et al., 2011).
@@
Our main results are tight asymptotic upper and lower bounds on the PoU for consensus games. Letting $n$ denote the number of agents participating in the consensus game, the previously best known asymptotic bounds due to Balcan et al. (2011) were a lower bound of $\Omega(n^{2}\varepsilon^{3})$ and an upper bound of $O(n^{2}\varepsilon)$. In this paper, we extend their work by proving that the PoU for consensus games is exactly $\Theta(n^{2}\varepsilon^{2})$. We believe that our result also holds irrespective of agent update ordering (i.e., the adversary can only control the magnitude of the perturbations), and this is a potential future direction for research.
@@ 2 The Model
We consider a consensus game in which each of $n$ agents is located at a distinct vertex of a simple undirected graph with $n$ vertices (Christodoulou et al., 2012). In this paper, we will use the words agent and vertex interchangeably. Each agent can choose a color from $\{0,1\}$. Let us refer to two agents as neighbors if and only if there exists an edge between them. An agent’s cost is defined as the number of its neighbors that pick a different color; we call these bad neighbors. We will also call edges whose endpoints have different colors bad edges, and we will call neighbors or edges that are not bad good. Finally, for any state $S$ of the game, we define a social cost function $\operatorname{cost}(S)$ as the sum of the costs of all agents in the game.
Starting from a ground state $S_{0}$ with social cost at least 1, let $S_{t}$ be a state achieved after $t$ steps under uncertain best-response dynamics. Here, the term uncertain refers to the fact that an agent might miscount its cost by a multiplicative factor of at most $1+\varepsilon$ for an uncertainty factor $0<\varepsilon<1$ that can depend on $n$. Define an uncertain best response as a best response that strictly increases the social cost as a result of uncertainty. In particular, the number of bad neighbors of the agent playing an uncertain best response must not exceed the number of good neighbors by more than a multiplicative factor of $1+\varepsilon$. Also, by definition, an uncertain best response must involve a color switch. Note that we can assume $\varepsilon=\Omega\left(\frac{1}{n}\right)$, otherwise an agent cannot miscount its number of bad neighbors and hence there is no uncertainty. The price of uncertainty (PoU) of consensus games is defined as
@@
This paper obtains a lower bound of $\Omega\left(\varepsilon^{2}n^{2}\right)$ on the PoU of consensus games for all $\varepsilon=\tilde{\Omega}(n^{-1/2})$. Here $\tilde{\Omega}$ hides factors that are logarithmic in $n$. This is achieved using an intricate construction that amplifies the “snowball” effect of uncertainty described in Balcan et al. (2011). Next, by analyzing a clever construction, we prove that the PoU for consensus games is also bounded from above by $O\left(\varepsilon^{2}n^{2}\right)$ for all $\varepsilon=\Omega(n^{-1/4})$. These bounds coincide, so we conclude that the asymptotically tight PoU for consensus games is $\Theta(\varepsilon^{2}n^{2})$ for $\varepsilon=\Omega(n^{-1/4})$, resolving an open problem of Balcan et al. (2011).
@@ 2.1 Main Technical Challenges
Lower Bound. In order to achieve a lower bound of $\Omega(\varepsilon^{3}n^{2})$, Balcan et al. (2011) utilizes an initializer gadget to set up initial colorings for an output component. This component consists of successive layers of vertices, each of which has more vertices than the previous layer by a factor of $(1+\varepsilon)$, and every vertex is connected to all vertices on adjacent layers. This output component is responsible for generating $\Omega(\varepsilon n^{2})$ bad edges, and the initializer gadget requires $\Theta(1/\varepsilon^{2})$ starting bad edges, so this achieves a lower bound of $\Omega(\varepsilon^{3}n^{2})$. We have found that it is very difficult to increase the final number of bad edges produced by such an output for any construction because the graph immediately becomes too connected for enough agents to play uncertain best responses. Rather, we present an alternative construction for an initializer gadget in Section 3.1 that only requires $\Theta(1/\varepsilon)$ starting bad edges, which results in a construction for which the number of bad edges can increase by a factor of $O(\varepsilon^{2}n^{2})$.
Upper Bound. Generally, the goal of an adversary that wants to blow up the number of bad edges can be thought of as trying to maximize the number of additional bad edges that each new vertex produces in the uncertain best response process. Consider a time in which there are $m^{2}$ bad edges. The most “compact” way to store these bad edges using the least number of vertices is a bipartite graph $K_{m,m}$. If each of the $m$ vertices on one side of the bipartite graph plays an uncertain best-response, we can blow up the number of bad edges by a multiplicative factor of $(1+\varepsilon)$. However, this requires $(1+\varepsilon)m$ new vertices to “absorb” the new bad edges produced by the uncertain best-response. Hence, increasing the number of bad edges from $m^{2}$ to $(1+\varepsilon)m^{2}$ requires roughly $(1+\varepsilon)m$ new vertices.
Following this heuristic for a constant $\varepsilon$ (i.e., not depending on $n$), the last stage will have a $K_{\Theta(\varepsilon n),\Theta(\varepsilon n)}$. Let us be optimistic in the sense that we assume that the original bad edges at each stage, which become good after vertices play an uncertain best response, can be turned bad again. Assuming most edges can eventually be turned bad again, as is the case in our lower bound construction, the total number of bad edges at the end is
@@
Since there must be $\Omega(1/\varepsilon)$ bad edges at the start, we estimate the upper bound on the price of uncertainty for consensus games to be $O(\varepsilon^{2}n^{2})$.
It appears natural to track the number of bad edges at every step and try to establish an upper bound on its growth rate as new vertices play uncertain best responses. The hope is that we can prove an upper bound on the growth rate in line with our estimate of $O(\varepsilon^{2}n^{2})$. Unfortunately, this approach fails, as illustrated by the following example.
@@
Consider a red vertex $V$ whose only neighbors are $m$ blue vertices and $m+1$ red vertices for $m=\Omega(1/\varepsilon)$. The $m$ blue vertices are all connected to another blue vertex $B$, and the $m+1$ red vertices are all connected to another red vertex $R$. Then, we can have $V$ play $2m+1$ uncertain best responses in order to roughly double the number of bad edges while only operating on these $O(m)$ vertices.
@@
  • Let $V$ play an uncertain best response that switches its color to blue, resulting in $m+1$ bad edges.
@@
  • Take the endpoint of one such bad edge adjacent to $R$ and switch its color so that there are now $m$ bad edges incident to $V$ and 1 bad edge incident to $R$.
@@
  • Switch the color of $V$ again, resulting in $m+1$ bad edges. Take the endpoint of one such bad edge adjacent to $B$ and switch its color so that there are now $m$ bad edges incident to $V$, 1 bad edge incident to $B$, and 1 bad edge incident to $R$.
@@
  • Repeat the above, switching at each step the color of a neighbor of $V$ whose color has not been switched before, until all neighbors of $V$ have had their colors switched once. There are now $m$ bad edges incident to $B$ and $m+1$ bad edges incident to $R$.
@@
Note that Example 1 can be generalized to operate on a $K_{m,2m+1}$ where each vertex in the $m$-size part is blue, $m$ of the vertices in the $(2m+1)$-size part are red, and $m+1$ of the vertices in the $(2m+1)$-size part are blue to also roughly double the number of bad edges. The bad edge multiplicative growth rate of this process operating on $\Theta(m)$ vertices is 2, whereas in our initial estimate it was $1+\varepsilon$. Hence, using this analysis, it is actually possible for there to be $\Omega(n^{2})$ bad edges at the last stage if operating on $m$ vertices can double the number of bad edges at each step. We thus cannot prove an upper bound of $O(\varepsilon n^{2})$ on the number of bad edges at the last stage using this strategy.
@@
Instead, one idea is that we should focus not on the total number of bad edges at every step but rather consider the step at which each vertex first plays an uncertain best response. This imposes a sequential order on the vertices from the first vertex to play an uncertain best response to the last. Recall that in order for a vertex with high degree to be able to play an uncertain best response, it must be the case that a large number of edges adjacent to this vertex must be bad. If we are to end up with $\Omega(\varepsilon n^{2})$ bad edges at the end, it must be the case that the later vertices in this sequence must have high degree on average. This requires a large number of existing bad edges to be generated from the vertices before the final vertices in the sequence, which in turn requires bad edges from the vertices before them, and so on.
@@
Our proof of the $O(\varepsilon^{2}n^{2})$ upper bound analyzes the degrees of vertices in this sequence and shows that the sequence of degrees cannot increase “too fast” if vertices can only play $(1+\varepsilon)$-uncertain best responses, where we realize the notion of “too fast” via a clever construction.
@@ 3 Lower Bound
We first improve the lower bound on the PoU for consensus games. Furthermore, we do so in the expanded domain $\varepsilon=\tilde{\Omega}(n^{-1/2})$. Here, $\tilde{\Omega}$ hides factors that are logarithmic in $n$.
@@ 3.1 Proof of Improved Lower Bound
One general idea is to construct a graph as a sequence of levels, where each level contains slightly more vertices than the previous level. At the beginning, all of the vertices in the first level should have enough bad edges to play uncertain best responses, which increases the number of bad edges by a multiplicative factor of $1+\varepsilon$. The fact that the vertices in the first level switched color implies that the vertices in the second level now have enough bad neighbors to play an uncertain best response, and this effect snowballs across the levels.
@@
We will make the above idea rigorous. We define a boosting operation between layers of the following form. Let $\ell>1$ and suppose that level $\ell-1$ has $\frac{1}{1+\frac{\varepsilon}{2}}\cdot m$ vertices for some $m$, level $\ell$ has $m$ vertices, and level $\ell+1$ has $\left(1+\frac{\varepsilon}{2}\right)\cdot m$ vertices. Suppose that every vertex in level $\ell$ is connected to every vertex in level $\ell-1$ and level $\ell+1$, so the degree of every vertex in level $\ell$ is $d_{\ell}=\left(\frac{1}{1+\frac{\varepsilon}{2}}+(1+\frac{\varepsilon}{2})\right)\cdot m$. If the vertices in levels $\ell$ and $\ell+1$ are all blue and the vertices in level $\ell-1$ are all red, then the edges between levels $\ell-1$ and $\ell$ are all bad. Since $(1+\varepsilon)\cdot\frac{1}{1+\frac{\varepsilon}{2}}\geq\frac{d_{\ell}}{2}$ for all $\varepsilon\leq 1$, then all vertices in level $\ell$ are allowed to play uncertain best responses in sequence. This turns level $\ell$ all red. Note now that the number of bad edges between levels $\ell$ and $\ell+1$ is $\left(1+\frac{\varepsilon}{2}\right)\cdot m^{2}$, which is a $1+O(\varepsilon)$ multiplicative increase from the number of original bad edges $\frac{1}{1+\frac{\varepsilon}{2}}\cdot m^{2}$ from before this boosting operation.
@@
We now present the construction. In order to implement the idea above, we must first find some way to create the first two layers of size $\frac{1}{\varepsilon}$ because a vertex must have degree at least $\frac{1}{\varepsilon}$ in order to play an uncertain best response. The initializer layers will have numbers of vertices
$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{1}{\varepsilon},\frac{1}{\varepsilon},$
@@
where the numbers of vertices in the odd-indexed layers cycle through the consecutive integers from $\frac{1}{\varepsilon}$ to $\frac{2}{\varepsilon}$, and the numbers of vertices in even-indexed layers increase by a factor of 2 whenever the odd-indexed layers’ cycle repeats up until the number reaches $\frac{1}{\varepsilon}$. Each vertex is connected to all vertices in the layer immediately after it except for vertices in odd-indexed layers with $\frac{2}{\varepsilon}$ vertices, which are connected to the next layer to form a regular bipartite graph with the next layer where the vertices in the next layer are each connected to $\frac{1}{\varepsilon}$ vertices in the current layer. Note that the total number of initializer layers is $\frac{1}{\varepsilon}\cdot\log_{2}\frac{1}{\varepsilon}$, and each layer has at most $\frac{2}{\varepsilon}$ vertices, so the total number of vertices in the initializer layers is $O\left(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon}\right)$. The initializer sequence has been constructed in such a way such that by starting from a state in which the first layer is red and all other layers are blue, the vertices in each layer can play uncertain best responses to change their color to red in sequence so that the last two layers form a $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ where the left side is red and the right half is blue. Finally, we construct $\frac{1}{\varepsilon}$ blue vertices in a parallel layer to the red left side of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$ and connect each of them to every vertex in the right half of the $K_{\frac{1}{\varepsilon},\frac{1}{\varepsilon}}$, so that each vertex in the right half now has $\frac{1}{\varepsilon}$ red neighbors and $\frac{1}{\varepsilon}$ blue neighbors on the left. Note that no new bad edges are added. We will now leave all layers before the last layer frozen: no vertex in a layer before the last layer will ever change color again.
@@
After the last layer in the initializer, construct a secondary set of layers with numbers of vertices
$\frac{1}{\varepsilon},1,\frac{1}{\varepsilon}+1,1,\ldots,\frac{2}{\varepsilon}-1,1,\frac{2}{\varepsilon},2,\frac{1}{\varepsilon},2,\frac{1}{\varepsilon}+1,\ldots,\frac{2}{\varepsilon}-1,2,\frac{2}{\varepsilon},4,\frac{1}{\varepsilon},\ldots,\frac{2}{\varepsilon},\frac{2}{\varepsilon}.$